Near Earth, the density of protons in the solar wind (a stream of particles from the Sun) is 9.66 cm-3, and their speed is 697 km/s. (a) Find the current density of these protons. (b) If Earth's magnetic field did not deflect the protons, what total current Earth receive?

(a) Number: _________________ Unit: A/m^2

(b) Number: _________________ Unit: A
Answer #1

Number density of protons = 9.66 (cm^(-3)) = 9.66e-6 (m^(-3))
Speed = 697(km/s) = 6.97e5 (m/s)

a) Current density:
J = n * q* v = 0.966e-5 * 6.97e5 * 1.6e-19 = = 10.77e-19 (C/(s*m^2)) = 10.77e-19 (A/m^2)

b) Total current:
I = J * cross-sectional area of Earth
= J * π * R_e^2
= 10.77e-19* π * (6371 km)^2
= 10.77*6.371^2*π * e(-7) = 1.37*e-4 (A)